a rope 82cm long is cut into two pieces with one peice 4 more than twice as long as the other. find the length of each piece.
Answer by 24HoursTutor.com(40)

Let the first piece of rope be x
The second piece of rope is then 4 + 2x
We know that together they were 82 cm. So we can write the equations as:
x+2x+4=82
3x+4=82
3x=82-4
3x=78
x=78/3
x=26, the length of the first piece of rope
4+2x=4+52=56, the length of the second piece of rope.
Ans.: The length of the two pieces of rope are 26cm and 56 cm.
The above question is solved by one of the experts from 24HoursTutor.com
We provide live 1:1 online homework help
1. 24X7 tutors available
2. Certified tutors
3. Only for $9.99/hr